4^2-4-4-4=8-4. Before I thought of moving the equation around. I am out of practice.....
Answer by jim_thompson5910(35256)

4^2 - 4 - 4 - 4 = 16 - 4 - 4 - 4

4^2 - 4 - 4 - 4 = 12 - 4 - 4

4^2 - 4 - 4 - 4 = 8 - 4

4^2 - 4 - 4 - 4 = 4

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------